OpenStudy (anonymous):

given that log(base 5)x=y, express log(base 25)x in terms of y

jimthompson5910 (jim_thompson5910):

\[\Large \log_{5}(x) = y\] \[\Large \frac{\log_{10}(x)}{\log_{10}(5)} = y\] \[\Large \log_{10}(x) = y\log_{10}(5)\] \[\Large 2\log_{10}(x) = 2y\log_{10}(5)\] \[\Large 2\log_{10}(x) = y\log_{10}(5^2)\] \[\Large 2\log_{10}(x) = y\log_{10}(25)\] \[\Large 2\frac{\log_{10}(x)}{\log_{10}(25)} = y\] \[\Large 2\log_{25}(x) = y\] \[\Large \log_{25}(x) = \frac{y}{2}\]

OpenStudy (anonymous):

or we could approach it this way, without changing the base \[\log_5(x)=y\iff 5^y=x\iff 25^{\frac{y}{2}}=x\iff \log_{25}(x)=\frac{y}{2}\]